Shift report 20020105 08:00-16:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Jan 05 2002 - 16:06:00 EST

  • Next message: Apache: "Shift report 20020106 16:00-24:00"

    ------------------------------------------------------------
    Date: Sat Jan  5 16:06:00 2002, Shift: 08:00-16:00
    Supervisor: Mads , with: FV
    Log book#12 pages 2 - 2
    
    Shift Summary:
    ------------------------------------------------------------
    There was only useable beam from 1510 to 1540 at which time Flemming was doing ajustments to some triggers. So there was no data run during this shift.
    
    Did pedestal runs 6283 and 6284.
    
    No pedestal runs for tpcs and DVM.
    No detector problems observed.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sat Jan 05 2002 - 16:06:48 EST